Opened 5 years ago

Closed 4 years ago

Last modified 3 years ago

#530 closed flight processing (fixed)

RG12/09, flight day 172/2014, Wessex

Reported by: knpa Owned by: benj
Priority: alpha 5 Milestone: 2014 data processing completion
Component: Archiving Keywords:
Cc: Other processors:

Description (last modified by tec)

Data location: /users/rsg/arsf/arsf_data/2014/flight_data/uk/RG12_09-2014_172_Wessex

Data arrived from ARSF via network transfer on 23/06/2014

RG12/09 Scientific objective: Unknown

RG12/09 Priority: a5

PI: France Gerard

Notes: There is OWL data. The last line is missing though, and the last 8 lines are missing blackbody data. There is FW data. Deleted 3 corrupted webcam imagesL 1020, 1015, 1037

Sensors:

Camera (Requested), Delivered 10/09/2014, Archvied 13/05/2015
Fenix (Requested), Delivered 10/09/2014, Archvied 13/05/2015
LiDAR (Requested), Delivered 08/10/2014, Archvied 13/05/2015
FW LiDAR (?), Delivered 08/10/2014, Archvied 13/05/2015

Change History (38)

comment:1 Changed 5 years ago by tec

Started navigation processing

HUNG Basestation Data
Latitude51 24 15.45169
Longitude-1 30 50.17037
El Height183.194m

comment:2 Changed 5 years ago by tec

Trimmed down the gps start and end times to 550800 - 560880.

Finished nav processing

comment:3 Changed 5 years ago by dap

Began Fenix processing.

comment:4 Changed 5 years ago by tec

Started LiDAR processing.

comment:5 Changed 5 years ago by dap

Beginning RCD processing

comment:6 Changed 5 years ago by dap

SCT values for Fenix data:

Flightline FENIX
1 0.95
2 0.97
3 0.99
4 0.98
5 0.95
6 0.90
7 0.90
8 0.90
9 0.97
10 0.93
11 0.95
12 1.10
13 1.09
14 0.96
15 0.90
16 0.90
17 0.93

Edit: Delivery check revealed that some of the SCT values were incorrect, so they have been changed in this comment.

Last edited 5 years ago by dap (previous) (diff)

comment:7 Changed 5 years ago by dap

Completed Fenix processing and created delivery and readme. Awaiting delivery check.

comment:8 Changed 5 years ago by dap

RCD processing complete, awaiting delivery check.

comment:9 Changed 5 years ago by tec

Starting RCD DC.

comment:10 Changed 5 years ago by tec

Found some issues:

  • In ReadMe where it displays a thumbnail for each RCD line, the lines are out of order, ie 10 comes after 18.
  • Photos 335-339, look a bit over exposed.

Thats it, RCD DC finished.

comment:11 Changed 5 years ago by dap

  • Asked mark1 about the ordering of the flight lines in the read me, this is the correct order since this is the order the lines were flown in.
  • Checked the images found to be over-exposed against Google Earth's satellite imagery and found that the camera images weren't over exposed.

RCD images ready for delivery.

comment:12 Changed 5 years ago by knpa

Beginning Fenix delivery check.

comment:13 Changed 5 years ago by knpa

Found a lot of problems with delivery so am going to fail this at this point and will check it properly when these are done. Be sure to check obvious things before you submit for delivery check such as none of the folders are empty.

Current issues:

  • No logsheet. Create and submit for checking.
  • ReadMe flightline filenames table is messed up. Easiest way to fix is to put /newpage before the table in the tex.
  • Remove flightlines/fodis directory as this is not needed with fenix.
  • No screenshots - need to create these
  • project_information/ dir is empty
  • No xsl or css files present in the doc/ directory (talk to mark1 or maybe stgo)

comment:14 Changed 5 years ago by dap

Fixed all problems found. Ready for re-delivery check.

comment:15 Changed 5 years ago by knpa

2nd Delivery Check of HS.

Major fail:
Looked at screenshots and found sct wobbles. Line 14 and Line 11 for example. Please go back and correct SCTS. Then resubmit for checking.

Also found minor points below:

  • Remove unzipped mapped files from directory
  • Screenshots - first file is missing
  • Screenshots - remove sct part from names
  • Would rewrite data quality remark slightly: "Some flightlines [or give lines if known] are partly obscured by cloud cover".
  • sensor_FOV_vectors/fenix_viewvector_list.txt should be a windows compatible text file

comment:16 Changed 5 years ago by dap

Correcting problems found in Hyperspectral

comment:17 Changed 5 years ago by dap

Changed SCT values to:

Flight line FENIX
11 0.95
14 0.96
4 0.98
5 0.95
Last edited 5 years ago by dap (previous) (diff)

comment:18 Changed 5 years ago by dap

Hyperspectral Fenix

Issues fixed, ready for delivery check.

comment:19 Changed 5 years ago by tec

LiDAR

Flightline Roll Pitch
090258 -0.005802596 -0.000857190
091241 -0.005802596 -0.000857190
092213 -0.006022596 -0.000857190
093055 -0.005402596 -0.000702190
094204 -0.006202596 -0.000857190
095055 -0.005202596 -0.000857190
100014 -0.005722596 -0.000857190
101015 -0.005722596 -0.000857190
101959 -0.005722596 -0.000857190
103315 -0.005722596 -0.000857190
105715 -0.005722596 -0.000857190
110516 -0.005722596 -0.000857190
111208 -0.005722596 -0.000857190
112040 -0.005722596 -0.000857190
112813 -0.005722596 -0.000857190
113544 -0.005722596 -0.000857190
114122 -0.005722596 -0.000857190
Last edited 5 years ago by tec (previous) (diff)

comment:20 Changed 5 years ago by knpa

Done Hyperspectral delivery check.

Minor problems:

  • Last band is blank. Check with Mark or Ben if this is supposed to be happening.
  • Should change Quality Issues bit in ReadMe that talks about cloudiness. I think it should say 'many' instead of 'some' - there are a LOT of clouds over this collection, as you can see in the mosaic screenshot
  • Make sure to zip up the mapped files before sending

comment:21 Changed 5 years ago by dap

Hyperspectral Fenix

I've spoken to benj about the last band being blank and he said this has been an issue before and that it's due to a problem with the calibration. This is, however, not a problem.

Also changed the data quality remarks and updated readme.

comment:22 Changed 5 years ago by dap

Copying RCD and Fenix data on to hard drive.

comment:23 Changed 5 years ago by dap

RCD and Fenix data delivered to Dr Gerard at Centre for Ecology and Hydrology, 10/09/2014.

comment:24 Changed 5 years ago by dap

  • Description modified (diff)

comment:25 Changed 5 years ago by dap

  • Component changed from Processing: general to Archiving
  • Owner set to benj

comment:26 Changed 5 years ago by tec

LiDAR
Finished creating a delivery, awaiting delivery check.

comment:27 Changed 5 years ago by lah

Lidar delivery check - Only minor issues

  • Why are there 19 flight lines, but only 17 in log? There are also only 17 .trj files.
  • Do we need an empty HSI page in log?
  • I have changed DEM header resolution to 2.0.
  • Mean height from demcompare is -0.34.
  • What are the actual roll and pitch offsets for wiki?
Version 0, edited 5 years ago by lah (next)

comment:28 Changed 5 years ago by tec

One line was split into 3 lines.
Nothing can be done for the logsheet, logsheet script really needs redoing.
Updated Pitch and Roll in wiki

comment:29 Changed 5 years ago by lah

Lidar Delivery Check Complete

I have removed the extra page in the logsheet, so ready to deliver. The logsheet script still needs changing.

Last edited 5 years ago by lah (previous) (diff)

comment:30 Changed 5 years ago by dap

  • Description modified (diff)

LiDAR data delivered to Dr Gerard, 08/10/2014

comment:31 Changed 4 years ago by tec

Archiving
Starting archiving

comment:32 Changed 4 years ago by tec

  • Description modified (diff)
  • Resolution set to fixed
  • Status changed from new to closed

Archiving
Done confirmed as of 13/05/2015 12:10

comment:33 Changed 4 years ago by lah

Fenix re-DC

  • Have removed empty Eagle & Hawk page from logsheet.
  • spectra now look good.
  • only 1 problem remains: there are a lot of overflows in the readme (all VNIR bands) which don't show up in fastQC. Suggest that we need to double check overflows and figure out what's going wrong (and how to fix fastQC) before delivery.

comment:34 Changed 4 years ago by dac

Fenix

Have checked radiance values against Py6S simulations and looks good. Zipping files and marking as ready to deliver.

comment:35 Changed 4 years ago by gej

Reprocessed Hyperspectral data sent off to Dr Gerrard on 21st August 2015 via USB Hard Disk

comment:36 Changed 4 years ago by asm

Archiving
Starting archiving due to reprocessing. Files currently being uploaded.

comment:37 Changed 4 years ago by asm

Archiving
Finished. Notification mail: on hold

Note: See TracTickets for help on using tickets.